OpenStudy (gudetamas):

Write an equation of the perpendicular bisector of the segment with the endpoints (8,10) and (−4,2).

OpenStudy (booksforlife):

Give me just a minute. I might be able to help. :)

OpenStudy (gudetamas):

ok thank you

OpenStudy (booksforlife):

Do you know how to find the midpoint?

OpenStudy (gudetamas):

yes y2-y1/x2-x1

OpenStudy (gudetamas):

so it would be 2-10/-4-8 = -8/-12 then it simplifies to -2/-3 and then you change it to the opposite reciprocal 3/2

OpenStudy (booksforlife):

That's the slope formula.

OpenStudy (booksforlife):

The midpoint is mp=(x1+x2/2, y1+y2/2)

OpenStudy (gudetamas):

oh then its x1+x2/2 and y1+y2/2

OpenStudy (gudetamas):

so -2+6/2 for x and 0+12/2 for y, so x=2 and y=6 ?

OpenStudy (booksforlife):

It would be 8+(-4)/2 and 10+2/2

OpenStudy (gudetamas):

ok yes sorry was looking at a different problem, so x=2 and y=6.

OpenStudy (booksforlife):

Yes! Now you will find the slope. Which you've already done.

OpenStudy (gudetamas):

yes than u put it into the formula y-y1=m(x-x1) so y-6=3/2(x-2) ?

OpenStudy (booksforlife):

You will want to put it in point-slope form. Which is y-k=m(x-h)

OpenStudy (gudetamas):

yes so y-6-3/2(x-2)

OpenStudy (booksforlife):

Looks right! Now solve the point-slope equation for y to get y=mx+b

OpenStudy (gudetamas):

so you multiply 3/2 and -2 so it equals to y-6=3/2x-3

OpenStudy (booksforlife):

You will want to distribute the slope value and move the k value to the right side of the equation.

OpenStudy (gudetamas):

ok so you plus 6 with -3 and it equals to 3 so y=3/2x+3?

OpenStudy (booksforlife):

Yes, I believe you are correct.

OpenStudy (danjs):

the original segment slope you had good to -2/-3, that is 2/3 simplified

OpenStudy (danjs):

the perpendicular slope is negative and flipped over -3/2

OpenStudy (danjs):

negative over negative is a + value.. .you can write -2/-3 as (-1*2)/(-1*3), which is (-1/-1) * (2/3), or 1*(2/3), positive value

OpenStudy (booksforlife):

If you get stuck again @gudetamas this website will help you. http://www.algebralab.org/studyaids/studyaid.aspx?file=Algebra2_10-4.xml And here is a video on how to work out problems like this. https://www.youtube.com/watch?v=sTu2rw1ib5o

OpenStudy (gudetamas):

ok, so then would you do -3/2 x -2/1 ?

OpenStudy (booksforlife):

Yes, I believe so.

OpenStudy (gudetamas):

do you the answer ?

OpenStudy (booksforlife):

Work it out with the video and see if you can get the answer. If you still need help tag me. :)

OpenStudy (gudetamas):

nvm its 9 i found the answer -3/2x+9

OpenStudy (booksforlife):

Great! Could I please get a medal?

OpenStudy (gudetamas):

ok but you didnt really help me alot though.

OpenStudy (mathmale):

@BooksForLife : Kindly do NOT ask for medals, or ask to be fan-ned. If you focus on doing a good job of leading someone else through a problem and finding his/her own answer, at least some users will automatically award you with a medal. Thank youl

OpenStudy (danjs):

The general thing, is you need 2 points, or 1 point and a direction , in order to graph a line neither are given, the point was the midpoint of a segment, the direction was perpendicular to that same segment

OpenStudy (gudetamas):

i know, i found the answer now thanks